OpenStudy (anonymous):

Solve for x:

OpenStudy (anonymous):

\[(7-4\sqrt{3})^{x^{2}-4x+3} + (7+4\sqrt{3})^{x^2-4x+3} = 14\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

hi firstly multiply it by \[(7+4\sqrt{3})^{x^2-4x+3}\]

OpenStudy (anonymous):

then let \[a=(7+4\sqrt{3})^{x^2-4x+3}\] to get a simple quadratic equation for a and solve for a

OpenStudy (anonymous):

just notice that \[(7-4\sqrt{3})(7+4\sqrt{3})=1\]

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

what about (7-4root3)^...................?

OpenStudy (anonymous):

it is not a simple quadratic eq.

OpenStudy (anonymous):

after multiplying it by (7+4root3)^................... u have \[((7+4\sqrt{3})(7-4\sqrt{3}))^{x^2-4x+3}+(7+4\sqrt{3})^{x^2-4x+3}(7+4\sqrt{3})^{x^2-4x+3}=14(7+4\sqrt{3})^{x^2-4x+3}\] use my hint to simplify the first term on the left hand side of the equation

OpenStudy (anonymous):

can u do that?

OpenStudy (anonymous):

@mukushla what eq. you got after replacing it by a.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[(7-4\sqrt{3})^{x^2-4x+3} + (7+ 4\sqrt{3})^{x^2-4x+3} = 14 \] lets start with @mukushla method... it looks good. \[(7+ 4\sqrt{3})^{x^2-4x+3} ((7-4\sqrt{3})^{x^2-4x+3} + (7+ 4\sqrt{3})^{x^2-4x+3}]\) = (7+ 4\sqrt{3})^{x^2-4x+3} (14) \]

ganeshie8 (ganeshie8):

\[(1)^{x^2-4x+3} + (7+ 4\sqrt{3})^{2(x^2-4x+3)}) = (7+ 4\sqrt{3})^{x^2-4x+3} (14) \] \[1+ (7+ 4\sqrt{3})^{2(x^2-4x+3)}) = (7+ 4\sqrt{3})^{x^2-4x+3} (14) \]

ganeshie8 (ganeshie8):

let a = \((7+ 4\sqrt{3})^{(x^2-4x+3)} \) --------------------(1) \(1 + a^2 = 14a\) -------------------------------------(2) we have a qudatratic equation now... u can try rest as below : 1) solve equation (2) for a 2) equate that with equation (1) 3) take log on both sides.. and solve the FINAL quadratic equation for the value of \(x\) Enjoy...

OpenStudy (anonymous):

@shubham.bagrecha sorry i lost my connection

OpenStudy (anonymous):

@ganeshie8 thank u